OpenStudy (anonymous):

@ganeshie8 At Hamilton University, it is observed that 90% of the students pass mathematics, 85% pass English, 78% pass both courses. What are the odds in favour of passing mathematics and failing English?

OpenStudy (anonymous):

 The Venn diagram is not in question i just made that to see it better, not sure if its right though.

OpenStudy (amistre64):



OpenStudy (amistre64):

the odds are 12 to 100-12

OpenStudy (anonymous):

Unfortunately i dont see how you got that

OpenStudy (anonymous):

but it is right

OpenStudy (amistre64):

assuming 100 people (pass math) - (pass math and pass english) = (pass math and fail english) agreed?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so 90-78 = 12 odds tend to be written as for:against such that for + against is equal to the total population count

OpenStudy (amistre64):

since we assumed 100: then 12 are for, and 100-12 are againt

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

or written another way: U = A + A' and odds are written as: A : A'

OpenStudy (anonymous):

Okay i see it now, thanks for the help :)

OpenStudy (amistre64):

good luck